OpenStudy (hannahmae187):

An alloy of tin is 15% tin and weighs 20 pounds. A second alloy is 10% tin. How many pounds of the second alloy must be added to the first to get a 12% mixture? A) 30lb. B) 40lb. C) 60lb.

OpenStudy (sshayer):

\[\frac{ 15 }{ 100 }\times 20+\frac{ 10 }{ 100 }\times x=\frac{ 12 }{ 100 }\times(20+x) \] x=?